Calculate the pH when 90.0 mL of 0.200 M HBr is mixed with 30.0 mL of 0.400 M CH₃NH₂ (Kb = 4.4 × 10⁻⁴).
Answer #1

CH3NH2 + HBr -------------> CH3NH3+ + Br-

initially

millimoles of CH3NH2 = 30.0 x 0.400 = 12.00

millimoles of HBr added = 90.0 x 0.20 = 18.0

excess HBr present .

excess HBr left = 18.0 - 12.0 = 6.0

total volume = 30 + 90 = 120 mL

[HBr] = 6.0 / 120 = 0.05 M

as HBr is strong acid

[HBr] = [H+] = 0.05 M

pH = - log [H+]

pH = - log [0.05]

pH = 1.30